OpenStudy (anonymous):

A pool is shaped like a cylinder, and has a volume of 125(pi) cubic feet. What is the volume of another pool has the same height but doubled diameter?

OpenStudy (anonymous):

say the height is 5 and the diameter is 10, the the volume would be \(\pi\times r^2h=\pi \times 5^2\times 5=125\pi\)

OpenStudy (anonymous):

since half the diameter is the radius

OpenStudy (anonymous):

double that and you would have a diameter of 20 making the radius 10 compute \[\pi\times 10^2\times 5\]

OpenStudy (anonymous):

we can do it without the numbers i chose if you like

OpenStudy (anonymous):

no i doubled the diameter from 10 to 20 so the radius got doubled from 5 to 10

OpenStudy (anonymous):

lets do it another way, without the made up numbers ok?

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

when you compute the volume it is the area of the circle times the height the area goes with the square, as in \(A=\pi r^2\) therefore, if you double the diameter i.e. multiply it by 2 the area of the circle is multiplied by \(2^2=4\) in other words the area will be 4 times as large making the volume also 4 times as large (since the height is fixed)

OpenStudy (anonymous):

so do I multiply \[125\pi \times4\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

which is the exact same thing you would get if you computed \[10^2\times 5\times \pi\]